# Boundary Effects and Oxygen Deficiency-Driven Pattern Transitions in Algal Bioconvection

### v1

@@ 3.1 Spiral-shaped boundaries induce large-scale rotational patterns
Figure 1: Representative top-view images of bioconvection patterns formed by CR cells confined within a spiral-shaped boundary with an open liquid-air interface. (A) Side-view schematic illustrating the experimental setup, showing a fluid depth of 5 mm, approximately 1 mm greater than the spiral boundary height. (B, D) Initial pattern formation at the fluid-air interface, where cells accumulate due to negative gravitaxis, forming lattice-like clusters. Over time, these clusters become increasingly disordered (F, H). Space–time plots (C, G, K) generated by stacking light intensity profiles along trajectories marked in panels B, F and J reveal pronounced rotational dynamics: inward-directed near the spiral core (K) and bidirectional rotation farther from the center (G). Fourier analysis (E, I) confirms a characteristic pattern wavelength of approximately 4–5 mm. See also Video 1 for the full dynamic evolution.
Figure 2: Additional representative images illustrating bioconvection pattern formation under similar experimental conditions to Fig. 1. Initially, cells accumulate at the fluid-air interface, forming distinct spiral-shaped regions of high cell density (B), which subsequently evolve into discrete, lattice-like structures. These structures appear as dense cell clusters (black in raw images, visualized as green spots in the post-processed image, panel C, separated by brighter, interconnected net-like regions of lower cell concentration (E). Space–time plots reveal inward-directed rotational motion near the spiral core (D), whereas areas farther from the center exhibit rotational motion in both clockwise and counterclockwise directions (G). Fourier analysis (I) confirms the characteristic wavelength of the bioconvection patterns to be approximately 5 mm. See also Video 2 for the full dynamic evolution.
@@
Two representative top-view of bioconvection patterns formed by CR cells in a system with a spiral-shaped boundary and an open liquid-air interface is shown in Figs. 1 and  2. In our experiments, the fluid depth is approximately 5 mm, roughly 1 mm greater than the height of the spiral boundary (Figs. 1A). Starting from a homogeneous suspension, cells swim upward due to negative gravitaxis, leading to a lower concentration of cells directly above the spiral boundary compared to the deeper surrounding areas. This results in a distinctive spiral-shaped cell accumulation at the fluid-air interface in areas between the spiral boundary (see Fig. 2B), which is susceptible to bioconvection instability. Consequently, this spiral arrangement fragments into smaller clusters, forming a lattice-like patterns characterized by distinct dark spots, which correspond to regions of high cell density (visualized as green in post-processed images). These dense clusters are separated by brighter, interconnected, net-like structures representing regions of lower cell density (appearing dark in post-processed images). Space-time plots generated along multiple trajectories across different regions of the setup reveal a pronounced large-scale rotational motion in the bioconvection patterns. Notably, near the center of the spiral, we dominantly observe a directed inward rotation toward the spiral core (Fig. 1K and Fig. 2D). In contrast, farther from the center, the rotational behavior becomes more complex, with bioconvection patterns exhibiting both clockwise and counterclockwise motion (Fig. 1G and Fig. 2G). Fourier analysis performed along various trajectories within the setup indicates that the characteristic wavelength of the patterns is approximately 4-5 mm measured in different parts of the channel (Fig. 1E, I and Fig. 2I).
@@
3.2 Oxygen deficiency trigger drastic pattern transitions in confined setups
@@
Figure 3: Evolution of bioconvection patterns under confined conditions with a closed top boundary. (A) Experimental setup with a transparent PMMA wall placed above the spiral structure, maintaining a 1 mm gap. (B–D) Initial lattice-like arrangement of high-density cell clusters (visualized as green spots in post-processed images) separated by lower-density regions. (F–H) Transition phase characterized by fragmentation and the formation of a disorganized, branched cellular network. (J–L) Reorganization into a new, stable pattern with a reduced wavelength. (E, I, M) One-dimensional Fourier transforms of light intensity profiles, extracted along the blue dashed lines in panels D, H, and L, reveal a reduction in the dominant wavelength from approximately 6 mm to 4 mm. See also supplementary Video 3 for the full dynamic evolution.
@@
We modified the boundary conditions by placing a solid transparent PMMA wall above the spiral structure, maintaining a 1 mm gap between the spiral wall and the top boundary. The PMMA-based chamber was completely filled with the cell suspension, ensuring the elimination of any air gaps or bubbles (Fig. 3A). Under these confined conditions, we observed a pronounced transition in the bioconvection patterns over time, which we attribute to oxygen depletion. To characterize this transition, we carried out frame-by-frame analysis, constructed detailed space-time plots, and applied one-dimensional Fourier transforms (1D-FT) to quantify changes in the pattern wavelengths. We measured the wavelengths at three critical stages: before the transition, at its onset, and after the transition. Long-duration recordings (approximately two hours) revealed a significant reduction in the characteristic wavelength, accompanying a dramatic reorganization of the bioconvection patterns.
@@
Similar to the setup with an open liquid-air interface, the bioconvection patterns initially emerged from a homogeneous suspension as localized accumulations of cells arranged in a lattice-like structure. These patterns were characterized by distinct dark regions, corresponding to areas of high cell density (visualized as green spots in post-processed images), separated by brighter regions of lower cell concentration (appearing dark in processed images) (Fig. 3B–C and Video 3). This lattice arrangement persisted for approximately 30–40 minutes before undergoing a drastic transition (Fig. 3F-M). The transition was marked by a temporary weakening or disappearance of the regular pattern, followed by the formation of a disorganized, branched network of cells at random locations (Fig. 3F–I), which rapidly percolated throughout the entire suspension, resulting in a brief intermediate phase of disorganization. Following this phase, the cells spontaneously reorganized into a new, stable, and more regular pattern with a smaller characteristic wavelength (Fig. 3J–M), which remained steady for over an hour as long as the cells retained their motility. Note that the upper portion of the transitional zone—highlighted by a red box in Fig. 3D—exhibits a temporal delay, revealing that the transition event propagates gradually across the suspension.To quantify the associated changes in the pattern wavelength, we applied one-dimensional Fourier transforms to the light intensity profiles extracted from the space-time plots at the selected time points. The results, presented in Fig. 3E, I, and M, reveal a clear reduction in the dominant wavelength from approximately 6 mm to 4 mm during the course of the transition.
@@
Figure 4: Bioconvection pattern evolution under fully confined conditions with no fluid gap. (A) Experimental setup where the top PMMA wall directly contacts the spiral-shaped boundary, eliminating the 1 mm fluid gap present in the previous configuration. (B–D) Sequential top-view images showing the disappearance or significant weakening of the initial bioconvection pattern, followed by reorganization into a new, stable pattern with a smaller characteristic wavelength. (E) Space–time plot illustrating the transition, highlighting two distinct regions with different stripe spacings, separated by a prolonged intermediate phase of disorganization. The longer transition time is attributed to reduced cell-to-cell communication caused by the absence of the fluid gap. Fourier analysis confirms a decrease in the dominant wavelength from approximately 5 mm to 3 mm (panel D); see also Video 4.
@@
Figure 5: Bioconvection pattern evolution under confined conditions with a 1 mm fluid gap (as sketched in Fig. 3A) in a PMMA-made setup containing circular obstacles. (A–C) Sequential top-view images showing the reorganization of bioconvection patterns, with a new pattern featuring a smaller characteristic wavelength emerging after approximately 50 minutes. (D–F) Corresponding processed images of the snapshots shown in panels A–C. (G-I) Space–time plots along the trajectories indicated in panels D–E, revealing the transition, characterized by two distinct regions with different stripe spacings; see also supplementary Video 5.
@@
Building on these observations, we conducted additional experiments under confined conditions, modifying the setup so that the top PMMA wall was in direct contact with the upper edge of the spiral-shaped boundary, thereby eliminating the 1 mm fluid gap present in the previous configuration (Fig. 4A). Sequential top-view images captured at different stages of bioconvection pattern evolution are shown in Fig. 4B–C. Under these fully confined conditions, the intermediate phase—during which the patterns disappeared or substantially weakened—was significantly prolonged, lasting approximately 50 minutes before the system reorganized into a new, stable configuration with a shorter characteristic wavelength (Fig. 4D). This prolonged transition and subsequent reorganization are clearly illustrated in the space–time plot (Fig. 4E), which reveals two distinct regions with different stripe spacings corresponding to pre- and post-transition states, separated by an extended intermediate phase characterized by weak or absent pattern formation. We attribute this delay to impaired fluid-mediated communication between cells due to the absence of the fluid gap, which in the previous configuration facilitated more efficient collective reorganization. Despite the slower dynamics, Fourier analysis confirms a reduction in the dominant pattern wavelength from approximately 5 mm to 3 mm following the transition. To further confirm that oxygen depletion is the underlying cause of this transition, we performed a control experiment in which the entire chamber, including the top boundary, was fabricated from air-permeable PDMS. Under otherwise identical conditions, no pattern transitions were observed (Fig. S1), strongly supporting the conclusion that restricted oxygen exchange in PMMA-based confined geometries is the key factor driving the observed drastic pattern reorganization.
@@ 3.3 Simulations
Figure 6: Top-view snapshots of cell density from 3D simulations of bioconvection patterns above the spiral-shaped boundary at $z=4$ mm. Either a zero-shear-stress condition (representing a free liquid-air interface) or a no-slip condition (representing a rigid top wall) is applied at the upper boundary (at $z=5$ mm). Initially, CR cells swim upward (negative gravitaxis) and accumulate near the surface, forming a distinct spiral-shaped cell density pattern (A, D). This structure becomes dynamically unstable, fragmenting into smaller, high-density clusters that evolve into down-welling plumes (see also Figs. 9-10). Notably, fragmented clusters, initially aligned parallel to the spiral walls (B), progressively reorient perpendicular to the boundaries over time (C, F). A comparison between panels B and E shows that fragmentation of the spiral-shaped accumulation occurs earlier in the system with the zero-shear-stress boundary condition. G) The space–time plot along the spiral trajectory in panel C, constructed from long-term simulation results, reveals the inward movement of plumes toward the spiral center, in good agreement with experimental observations. The color map has been slightly adjusted to enhance visualization; see also Videos 6 and 7.
@@
Figure 7: Cell density distribution at the bottom of the setup ($z=0$) from 3D simulations of bioconvection patterns. Snapshots at $t=100$ s reveal that the cell concentration at the bottom is approximately half that near the surface (see Fig. 6), and the cell clusters remain more diffuse and unlike the near-surface structures, clusters at the bottom do not exhibit reorientation perpendicular to the spiral boundaries, highlighting the depth-dependent development of bioconvection patterns; see Videos 8 and 9 .
@@
Figure 6 shows a representative top-view of bioconvection patterns that form directly above the spiral-shaped boundary at a height of $z=4$ mm, where $z=0$ corresponds to the bottom of the chamber and $z=5$ mm marks the liquid–air interface. At early times, CR cells exhibit negative gravitaxis, swimming upward against gravity and accumulating near the free liquid-air interface (see also the side view in Fig. 9 and the three-dimensional density profile in Fig. 8A). Owing to the shallower fluid layer directly above the spiral walls—approximately 1 mm compared to 5 mm in surrounding regions—the cell concentration at the surface is initially lower above the spiral boundaries than in adjacent areas. This variation in fluid depth seeds the formation of a distinct spiral-shaped accumulation pattern between the spiral walls, as shown in Fig. 6A, D and Fig. S2A, D, in agreement with experimental observations (Fig. 2B). However, this spiral-shaped accumulation is dynamically unstable due to density inversion, as CR cells are approximately 5% denser than water. As a result, the initially continuous spiral structure fragments into smaller, high-density clusters (Figs. 6B, E), which subsequently evolve into distinct downwelling plume structures (Fig. 9). The formation of these downwelling plumes is accompanied by the emergence of surface flows, as illustrated in Fig. 8B. Notably, the fragmented clusters at $z=4$ mm, which initially align parallel to the spiral walls, progressively reorient perpendicular to the boundaries over time, as depicted in Figs. 6C and F.
@@
Figure 8: (A) Three-dimensional view of the developing bioconvection patterns inside the setup. (B) Zoomed-in top view of the cell density color map and surface flow field with streamlines at $z=4$ mm, highlighting the formation of downwelling plumes. (C) Vertical cut-plane used to extract the density profiles and flow fields shown in Figs. 910; see also Video 10.
@@
Figure 9: Cell density distribution and corresponding flow field with streamlines extracted along a vertical cut plane passing through the center of the spiral-shaped boundary (as defined in Fig. 8C). Panel B shows a zoomed-in view of the region highlighted by the dashed red box in panel A, revealing the formation of large-scale vortices and strong downward plumes, with fluid velocities reaching approximately 1 mm/s—nearly an order of magnitude greater than the typical swimming speed of individual CR cells. The dashed orange box in panel A indicates the region further magnified in Fig. 10; see Video 11.
Figure 10: Cell density distribution and corresponding iso-density contour lines during the formation of downwelling plumes, showing zoomed-in views of the region highlighted by the orange box in Fig. 9 at successive time points.
@@
This reorganization of cell clusters near the surface raises the question of how the underlying bioconvection structures develop throughout the depth of the suspension. To investigate this, we examine the cell density distribution at the bottom of the setup. Figure 7 shows the cell density color map at $z=0$. Interestingly, comparing the snapshots at $t=100$ sec with the corresponding near-surface patterns in Fig. 6 reveals that not only the cell concentration is approximately two times lower at the bottom, but also that the cell clusters remain more diffuse and are not oriented perpendicular to the spiral boundary, in contrast to the reorganization observed near the surface in Figs. 6C and F.
@@
In addition to characterizing the evolving cell density patterns, we investigated the large-scale coherent flow fields generated as downwelling plumes developed within the suspension. To quantify the fluid dynamics associated with bioconvection, velocity fields were extracted along a vertical cut-plane intersecting the center of the spiral-shaped boundary, as defined in Fig. 8C. Figure 9 presents the cell density distribution together with the corresponding velocity field and overlaid streamlines in this plane. To further illustrate the onset of flow instabilities, Fig. 10 displays iso-density contour lines highlighting the emergence and structure of the developing plumes. Consistent with previous observations reported in the literature, we find the formation of vortices featuring strong downward flows concentrated within the plumes, with velocities reaching approximately 1 mm/s—nearly an order of magnitude greater than the typical swimming speed of individual CR cells. These results highlight how collective cell motion dramatically enhances fluid transport far beyond the capabilities of isolated swimmers.

### v3

@@ 3.1 Spiral-shaped boundaries induce large-scale rotational patterns
Fig.  1: (A-K) Representative top-view images of bioconvection patterns formed by CR cells confined within a spiral-shaped boundary with an open liquid–air interface show that, across nearly the entire setup, the patterns exhibit dominantly directed movement toward the spiral’s center. Panel (A) presents a side-view schematic of the experimental setup, indicating a fluid depth of 5 mm, approximately 1 mm higher than the height of the spiral boundary. Panels (B) and (D) show the early stages of pattern formation at the liquid–air interface, where cells accumulate due to negative gravitaxis, forming lattice-like clusters. As time progresses, these clusters become disordered, as seen in panels (F) and (H). Space–time plots in panels (C), (G), and (K), generated by stacking light intensity profiles along trajectories marked in panels (B), (F), and (J), reveal prominent inward-directed rotational dynamics across much of the domain. One-dimensional Fourier analysis along the dashed blue lines in panels (C) and (G) confirms a dominant pattern wavelength of approximately 4–5 mm, as shown in panels (E) and (I); see Video 1. (L-S) An additional representative experiment under the same conditions as panel (A), except that pattern movement is inward near the spiral’s center (M) but proceeds either clockwise (CW) or counterclockwise (CCW) farther out (Q). Initially, cells accumulate at the liquid–air interface to form distinct spiral-shaped high-density regions (N), which then evolve into discrete, lattice-like structures (L, O, P and R); see Video 2.
@@
Two representative top-view examples of bioconvection patterns formed by CR cells in a system with a spiral-shaped boundary and an open liquid–air interface are shown in Fig. 1. In these experiments, the fluid depth is approximately 5 mm, about 1 mm higher than the height of the spiral boundary (Fig. 1A). Starting from a homogeneous suspension, the cells swim upward due to negative gravitaxis, leading to a lower cell concentration directly above the spiral boundary relative to the deeper surrounding regions. This results in a characteristic spiral-shaped accumulation of cells at the liquid–air interface in the regions between the spiral walls (Fig. 1N), which becomes susceptible to bioconvection instabilities. As the system evolves, this initial configuration fragments into smaller clusters, forming lattice-like patterns characterized by distinct dark spots—indicative of regions with high cell density—rendered as green in the post-processed images. These dense clusters are separated by brighter regions corresponding to areas of lower cell density, which appear dark in the post-processed images. Space–time plots generated along multiple trajectories across the setup reveal prominent large-scale rotational motion in the bioconvection patterns. Figure 1A-K presents a representative experiment (observed in approximately 10% of our total $N\approx 50$ experiments) in which the patterns exhibit dominant inward movement toward the spiral center across nearly the entire setup (Fig. 1C, G, K). In contrast, Fig. 1L-S illustrates the more commonly observed scenario, where directed inward rotational motion is primarily seen near the spiral core (Fig. 1M), while regions farther from the center exhibit both clockwise (CW) and counterclockwise (CCW) rotational dynamics, with each region typically maintaining its initial direction of rotation over time (Fig. 1Q). Fourier analysis performed along the trajectories marked by blue dashed lines in Fig. 1C, G and Q reveals a characteristic pattern wavelength of approximately 4–5 mm, consistently observed across various regions of the setup (Fig. 1E, I, S).
@@
3.2 Oxygen deficiency triggers drastic pattern transitions in air impermeable setups
@@
Fig.  2: Evolution of bioconvection patterns under confined air-impermeable conditions with a closed top boundary. (A) Experimental setup with a transparent PMMA wall placed above the spiral structure, maintaining a 1 mm gap. (B–D) Initial lattice-like arrangement of high-density cell clusters (visualized as green spots in post-processed images) separated by lower-density regions. (F–H) Transition phase characterized by fragmentation and the formation of a disorganized, branched cellular network. (J–L) Reorganization into a new, stable pattern with a reduced wavelength. (E, I, M) One-dimensional Fourier transforms of light intensity profiles, extracted along the blue dashed lines in panels D, H, and L, reveal a reduction in the dominant wavelength from approximately 6 mm to 4 mm. See also supplementary Video 3 for the full dynamic evolution.
@@
We modified the boundary conditions by placing a solid transparent PMMA wall above the spiral structure, maintaining a 1 mm gap between the spiral wall and the top boundary. The PMMA-based chamber was completely filled with the cell suspension, ensuring the elimination of any air gaps or bubbles (Fig. 2A). Under these confined, air-impermeable conditions, we observed a pronounced transition in the bioconvection patterns over time, which we attribute to oxygen depletion. To characterize this transition, we carried out frame-by-frame analysis, constructed detailed space-time plots, and applied one-dimensional Fourier transforms (1D-FT) to quantify changes in the pattern wavelengths. We measured the wavelengths at three critical stages: before the transition, at its onset, and after the transition. Long-duration recordings (approximately two hours) revealed a significant reduction in the characteristic wavelength, accompanying a dramatic reorganization of the bioconvection patterns.
@@
Similar to the setup with an open liquid-air interface, the bioconvection patterns initially emerged from a homogeneous suspension as localized accumulations of cells arranged in a lattice-like structure. These patterns were characterized by distinct dark regions, corresponding to areas of high cell density (visualized as green spots in post-processed images), separated by brighter regions of lower cell concentration (appearing dark in processed images); see Fig. 2B–C and Video 3. This lattice arrangement persisted for approximately 30–40 minutes before undergoing a drastic transition (Fig. 2F-M). The transition was marked by a temporary weakening or disappearance of the regular pattern, followed by the formation of a disorganized, branched network of cells at random locations (Fig. 2F–I), which rapidly percolated throughout the entire suspension, resulting in a brief intermediate phase of disorganization. Following this phase, the cells spontaneously reorganized into a new, stable, and more regular pattern with a smaller characteristic wavelength (Fig. 2J–M), which remained steady for over an hour as long as the cells retained their motility. Note that the upper portion of the transitional zone—highlighted by a red box in Fig. 2D—exhibits a temporal delay, revealing that the transition event propagates gradually across the suspension. To quantify the associated changes in the pattern wavelength, we applied one-dimensional Fourier transforms to the light intensity profiles extracted from the space-time plots at the selected time points. The results presented in Fig. 2E and M reveal a reduction in the dominant wavelength from roughly $6\,\mathrm{mm}$ to $4\,\mathrm{mm}$ during the transition. This $\sim 2\,\mathrm{mm}$ decrease was consistent across all experiments exhibiting a pattern transition.
@@
Fig.  3: Bioconvection pattern evolution under fully confined air-impermeable conditions with no fluid gap. (A) Experimental setup where the top PMMA wall directly contacts the spiral-shaped boundary, eliminating the 1 mm fluid gap present in the previous configuration. (B–D) Sequential top-view images showing the disappearance or significant weakening of the initial bioconvection pattern, followed by reorganization into a new, stable pattern with a smaller characteristic wavelength. (E) Space–time plot illustrating the transition, highlighting two distinct regions with different stripe spacings, separated by a prolonged intermediate phase of disorganization. The extended transition time is attributed to weakened cell-to-cell hydrodynamic interactions due to the absence of a fluid gap. Fourier analysis confirms a decrease in the dominant wavelength from approximately 5 mm to 3 mm (panel D); see Video 4 for the full dynamic evolution.
@@
Fig.  4: Bioconvection pattern evolution under confined air-impermeable conditions with a 1 mm fluid gap (as sketched in Fig. 2A) in a PMMA-made setup containing circular obstacles. (A–C) Sequential top-view images showing the reorganization of bioconvection patterns, with a new pattern featuring a smaller characteristic wavelength emerging after approximately 50 minutes. (D–F) Corresponding processed images of the snapshots shown in panels A–C. (G-I) Space–time plots along the trajectories indicated in panels D–F, revealing the transition, characterized by two distinct regions with different stripe spacings; see Video 5.
@@
Building on these observations, we conducted additional experiments under confined conditions, modifying the setup so that the top PMMA wall was in direct contact with the upper edge of the spiral-shaped boundary, thereby eliminating the 1 mm fluid gap present in the previous configuration (Fig. 3A). Sequential top-view images captured at different stages of bioconvection pattern evolution are shown in Fig. 3B–C. Under these fully confined air-impermeable conditions, the intermediate phase—during which the patterns disappeared or substantially weakened—was significantly prolonged, lasting approximately 50 minutes before the system reorganized into a new, stable configuration with a shorter characteristic wavelength (Fig. 3D). This prolonged transition and subsequent reorganization are clearly illustrated in the space–time plot (Fig. 3E), which reveals two distinct regions with different stripe spacings corresponding to pre- and post-transition states, separated by an extended intermediate phase characterized by weak or absent pattern formation. We attribute this delay to reduced fluid-mediated hydrodynamic interaction between cells due to the absence of the 1 mm fluid gap, which in the previous configuration facilitated more efficient collective reorganization. Despite the slower dynamics, Fourier analysis confirms a reduction in the dominant pattern wavelength from approximately 5 mm to 3 mm following the transition. To further confirm that oxygen depletion is the underlying cause of this transition, we performed a control experiment in which the entire chamber, including the top boundary, was fabricated from air-permeable PDMS. Under otherwise identical conditions, no pattern transitions were observed (Fig. S1), strongly supporting the conclusion that restricted oxygen exchange in PMMA-based confined geometries is the key factor driving the observed drastic pattern reorganization.
@@ 3.3 Simulations
Fig.  5: Top-view snapshots of cell density from 3D simulations of bioconvection patterns above the spiral-shaped boundary at $z=4$ mm. Either a zero-shear-stress condition (representing a free liquid-air interface) or a no-slip condition (representing a rigid top wall) is applied at the upper boundary (at $z=5$ mm). Initially, CR cells swim upward (negative gravitaxis) and accumulate near the surface, forming a distinct spiral-shaped cell density pattern (A, D). This structure becomes dynamically unstable, fragmenting into smaller, high-density clusters that evolve into down-welling plumes (see also Fig. 7). Notably, fragmented clusters, initially aligned parallel to the spiral walls (B), progressively reorient perpendicular to the boundaries over time (C, F). A comparison between panels B and E shows that fragmentation of the spiral-shaped accumulation occurs earlier in the system with the zero-shear-stress boundary condition. G) The space–time plot along the spiral trajectory in panel C, constructed from long-term simulation results, reveals the inward movement of plumes toward the spiral center, in good agreement with experimental observations. The color map has been slightly adjusted to enhance visualization; see also Videos 6-8.
@@
Figure 5 shows a representative top-view of bioconvection patterns that form directly above the spiral-shaped boundary at a height of $z=4$ mm, where $z=0$ corresponds to the bottom of the chamber and $z=5$ mm marks the liquid–air interface. At early times, CR cells exhibit negative gravitaxis, swimming upward against gravity and accumulating near the free liquid-air interface (see also the side view in Fig. 7 and the three-dimensional density profile in Fig. 6A). Owing to the shallower fluid layer directly above the spiral walls—approximately 1 mm compared to 5 mm in surrounding regions—the cell concentration at the surface is initially lower above the spiral boundaries than in adjacent areas. This variation in fluid depth seeds the formation of a distinct spiral-shaped accumulation pattern between the spiral walls, as shown in Fig. 5A, D and Fig. S3A, D, in agreement with experimental observations (Fig. 1N). However, this spiral-shaped accumulation is dynamically unstable due to density inversion, as CR cells are approximately 5% denser than water. As a result, the initially continuous spiral structure fragments into smaller, high-density clusters (Figs. 5B, E), which subsequently evolve into distinct downwelling plume structures (Fig. 7). The formation of these downwelling plumes is accompanied by the emergence of surface flows, as illustrated in Fig. 6B. Notably, the fragmented clusters at $z=4$ mm, which initially align parallel to the spiral walls, progressively reorient perpendicular to the boundaries over time, as depicted in Figs. 5C and F.
@@
Fig.  6: (A) Three-dimensional visualization of emerging bioconvection patterns within the setup, represented by iso-surface concentration contours. (B) Zoomed-in top view of the cell density color map and surface flow field with streamlines at $z=4$ mm, highlighting the formation of downwelling plumes. (C) Vertical cut-plane used to extract the density profiles and flow fields shown in Fig. 7; see also Video 11 and 12.
@@
Fig.  7: Cell density distribution and corresponding flow field with streamlines extracted along a vertical cut plane passing through the center of the spiral-shaped boundary (as defined in Fig. 6C). Panel B shows a zoomed-in view of the region highlighted by the dashed red box in panel A, revealing the formation of large-scale vortices and strong downward plumes, with fluid velocities reaching approximately 1 mm/s—nearly an order of magnitude greater than the typical swimming speed of individual CR cells. The dashed orange box in panel A indicates the region further magnified in panel C. (C) Cell density distribution and corresponding iso-density contour lines during the formation of downwelling plumes, showing zoomed-in views of the region highlighted by the orange box in panel A at successive time points; see Videos 13 and 14.
@@
This reorganization of cell clusters near the surface raises the question of how the underlying bioconvection structures develop throughout the depth of the suspension. To investigate this, we examine the cell density distribution at the bottom of the setup. Figure S2 shows the cell density color map at $z=0$. Interestingly, comparing the snapshots at $t=100$ sec with the corresponding near-surface patterns in Fig. 5 reveals that not only the cell concentration is approximately two times lower at the bottom, but also that the cell clusters remain more diffuse and are not oriented perpendicular to the spiral boundary, in contrast to the reorganization observed near the surface in Figs. 5C and F.
@@
In addition to characterizing the evolving cell density patterns, we investigated the large-scale coherent flow fields generated as downwelling plumes developed within the suspension. To quantify the fluid dynamics associated with bioconvection, velocity fields were extracted along a vertical cut-plane intersecting the center of the spiral-shaped boundary, as defined in Fig. 6C. Figure 7 presents the cell density distribution together with the corresponding velocity field and overlaid streamlines in this plane. To further illustrate the onset of flow instabilities, Fig. 7C displays iso-density contour lines highlighting the emergence and structure of the developing plumes. Consistent with previous observations reported in the literature 14, we find the formation of vortices featuring strong downward flows concentrated within the plumes, with velocities reaching approximately 1 mm/s—nearly an order of magnitude greater than the typical swimming speed of individual CR cells. These results highlight how collective cell motion dramatically enhances fluid transport far beyond the capabilities of isolated swimmers.